if x is normally distributed with mean 12 and standard deviation 4 find p(x>=20)
1
Expert's answer
2021-08-29T17:11:34-0400

Solution:

"\mu=12,\sigma=4
\\X\sim N(\mu,\sigma)"

"P(X\ge20)=1-P(X<20)
\\=1-P(z<\dfrac{20-12}{4})
\\=1-P(z<2)
\\=1-0.97725
\\=0.02275"